OpenStudy (anonymous):

A boat A travels due west at a speed of 30km/h. The velocity of a boat B relative to A is 14km/h due south. Find the speed of boat B and the direction in which it is moving.

OpenStudy (dumbcow):

 use pythagorean thm to find x x^2 = 30^2 + 14^2 tan(theta) = 14/30